# Interactive-KBQA: Multi-Turn Interactions for Knowledge Base Question Answering with Large Language Models

### v1

@@ Abstract
This study explores the realm of knowledge-base question answering (KBQA). KBQA is considered a challenging task, particularly in parsing intricate questions into executable logical forms. Traditional semantic parsing (SP)-based methods require extensive data annotations, which result in significant costs. Recently, the advent of few-shot in-context learning, powered by large language models (LLMs), has showcased promising capabilities. Yet, fully leveraging LLMs to parse questions into logical forms in low-resource scenarios poses a substantial challenge. To tackle these hurdles, we introduce Interactive-KBQA, a framework designed to generate logical forms through direct interaction with knowledge bases (KBs). Within this framework, we have developed three generic APIs for KB interaction. For each category of complex question, we devised exemplars to guide LLMs through the reasoning processes. Our method achieves competitive results on the WebQuestionsSP, ComplexWebQuestions, KQA Pro, and MetaQA datasets with a minimal number of examples (shots). Importantly, our approach supports manual intervention, allowing for the iterative refinement of LLM outputs. By annotating a dataset with step-wise reasoning processes, we showcase our model’s adaptability and highlight its potential for contributing significant enhancements to the field.
@@
Guanming Xiong1, Junwei Bao, Wen Zhao1 1Peking University, Beijing, China {gm_xiong,zhaowen}@pku.edu.cn baojunwei001@gmail.com
@@ 1 Introduction
Knowledge-base question answering (KBQA) is an increasingly significant research area that leverages structured knowledge bases (KBs) to provide precise answers to natural language (NL) questions. A KBQA system aims to harness the vast, structured information residing in KBs, such as Freebase (Bollacker et al., 2008), Wikidata (Vrandečić and Krötzsch, 2014), or domain-specific databases like the Movie KB (Zhang et al., 2018). Accurately interpreting and responding to user inquiries with data from knowledge bases has potent applications across numerous sectors, making it a focal point of both academic research and industrial innovation.
@@
Underutilization of large language models (LLMs). Despite the demonstrated capabilities of LLMs in reasoning and few-shot learning, existing KBQA approaches have not fully leveraged these strengths. Most current methods either use LLMs as classifiers to identify predicates (Gu et al., 2023; Sun et al., 2024) or as mimics to generate possible logical forms or questions (Li et al., 2023a, b). There remains a significant opportunity to more effectively leverage LLMs in a few-shot setting, to enhance the accuracy and interpretability of KBQA systems.
To address these challenges, we introduce the Interactive-KBQA framework, which combines the reasoning power of large language models (LLMs) with three tools to interact with knowledge bases (KBs). By conceptualizing the LLM as an agent and the KB as the environment, Interactive-KBQA facilitates an iterative, dialogue-based problem-solving process. Figure 1 illustrates the overall process. The process adheres to the thought-action paradigm. Given a complex query, the LLM is required to think and provide an action to interact with the KB through a set of specific tools. These tools return the execution results as observations. Specifically, this paper designs a set of tools that support heterogeneous databases (DBs) (i.e., Freebase, Wikidata, and a Movie KB) with a unified interaction logic. We categorize complex questions and annotated only two exemplars with complete interactive processes for each type as in-context learning demonstrations to prompt the LLM to complete the task. Furthermore, the method introduced in this study allows for manual intervention. Consequently, we annotated a small dataset by hand, providing a detailed reasoning process, thereby creating a low-resource dataset. Finally, we fine-tuned open-source LLMs on this dataset. The experiments conducted have shown that this method is effective in low-resource contexts. We will release this high-quality dataset to further contribute to NLP research in the community.
@@
  • Design a unified SPARQL-based toolset and interaction logic that efficiently addresses a wide array of complex queries.
@@ 2 Related Work
Knowledge-base question answering (KBQA) methods can be broadly classified into two distinct categories: information retrieval (IR)-based methods and semantic parsing (SP)-based methods. These approaches address the challenge of mapping natural language (NL) queries to the structured formats of knowledge bases (KBs) effectively.
@@
SP-based Methods parse questions into executable logical forms and perform queries against the KB to retrieve answers. Initial works translated questions into intermediate logical forms before execution. (Yih et al., 2015) defined a query graph that resembles subgraphs of the KB and can be directly mapped to a logical form. (Hu et al., 2018; Lan and Jiang, 2020) extended query graphs to include aggregation operators, and (Chen et al., 2021b) enhanced candidate query generation by leveraging predictions of query structure.
In recent years, generative models have increasingly been used to directly generate logical forms. (Chen et al., 2021a) designed a retriever to fetch relevant KB items and utilized a Grammar-based Decoder based on LSTM for generating S-expressions (Gu et al., 2021). Meanwhile, (Das et al., 2021) employed case-based reasoning for the KBQA task, using Big Bird (Zaheer et al., 2020) to generate logical forms by retrieving relevant cases. Notably, works such as (Ye et al., 2022), (Shu et al., 2022), (Zhang et al., 2023), and (Yu et al., 2023) involved initially retrieving basic elements from the KB, such as entities, relations, subgraphs, or texts as supplementary information, before directly generating logical forms using the T5 model (Raffel et al., 2020). However, this approach requires substantial amounts of training data and suffers from a lack of transparency in the reasoning process.
Large language models (LLMs) for KBQA. Leveraging the potent few-shot capabilities, recent studies have utilized LLMs to improve reasoning over KBs. In the context of few-shot scenarios, Gu et al. (2023) proposed that LLMs should prioritize evaluating the plausibility of agent plans over directly generating answers. Concurrently, Li et al. (2023a) advocates for the generation of logical forms as preliminary drafts, with the subsequent refinement of these drafts into executable queries by relying on the knowledge base. To address multi-hop questions, Jiang et al. (2023b) has developed specialized interfaces for accessing the KB, whereas Sun et al. (2024) introduces a novel approach enabling LLMs to iteratively employ beam search reasoning on a knowledge base.
However, despite employing LLMs, these methods lack the exploration of more complex questions.
@@ 3.1 Problem Formulation
This study investigates a semantic parsing (SP) method for knowledge-based question answering (KBQA). A knowledge base (KB) can be formally represented as $K\in E\times R\times(E\cup L\cup C)$, where $E$ represents the set of entities, $R$ denotes the set of relations between entities, and $C$ denotes the set of classes. Given a question $Q$ and knowledge base $\mathcal{K}$, our goal is to generate an executable SPARQL expression $S$ that is consistent with the question. Given the above definitions, the task can be formalized as $p(S|Q,\mathcal{K})$.
@@ 3.2 Overview
Recent advancements in large language models (LLMs) have demonstrated remarkable abilities in few-shot learning and reasoning capabilities. However, fully leveraging LLMs to address complex KBQA challenges remains a challenge. To bridge this gap, we introduce Interactive-KBQA, an interactive method for KBQA that conceptualizes the LLM as an agent and the KB as an environment. This approach enables semantic parsing and SPARQL generation through dialogic interactions. Specifically, we design a unified interaction logic with three generic tools, capable of supporting various complex types of questions (e.g., type constraints, count-based queries, etc.) across multiple databases (e.g., Freebase, Wikidata, Movie KB). Figure 2 presents an illustrative example of the interactive process.
@@ 3.3 Tools for Knowledge Base
Design Principles of Tools SP-based KBQA method requires locating the element and identifying the appropriate graph pattern. Therefore, it is imperative to design tools at an atomic level to ensure their universality. Adhering to this principle, we introduce the following three tools:
SearchNodes(name): This function searches for nodes in the KB using the given surface name name. Its primary function is entity linking. Thus, it returns not only the formal name of a node in the KB but also provides distinguishing features of the entity, such as its description and type. Importantly, this tool bypasses state-of-the-art Entity Linking (EL) techniques in favor of a generic retrieval approach. Further details are listed in Appendix A.2.
SearchGraphPatterns(sparql, semantic): This function aims to identify and rank graph predicates within the KB that are essential, guided by the semantic parameter. The function requires the input sparql, which should consist of a SPARQL query beginning with "SELECT ?e WHERE ". Following this, it performs a query on the one-hop subgraph centered around the entity ?e. This includes both incoming and outgoing edges. Subsequently, it ranks the retrieved triples based on the semantic relevance to the semantic parameter and the predicates in triples. Ultimately, the tool returns the top K triples. This tool is designed to precisely identify subgraphs while discarding irrelevant information, thereby optimizing the context window’s use in LLMs. It supports flexible operations and is specifically optimized for Compound Value Type [^3] A Compound Value Type (CVT) is a Type within Freebase designed to represent data where each entry is composed of multiple fields. (CVT) in Freebase. For example, to find movies featuring Tom Hanks in Freebase, the usage will be: SearchGraphPatterns(’SELECT ?e WHERE {?e type.object.name "Tom Hanks"@en .}’, semantic="play in film"), and return: [(?e, film.actor.film -> film.performance.film, "Nothing in Common"), ...]. More usages are described in Appendix 3.
@@ 3.4 Interactive Process
Given a question $Q$, we first construct a prompt text
@@
where the instruction text Inst consists of tool descriptions, tool usage, and the format of interaction. Inst is database (DB) specific and pre-written. $E$ denotes a set of exemplars, and for each type of question, we manually annotated two complete examples in an interactive format.
In each turn $T$, we let the LLM generate a action based on the Prompt and the history $H$ of the interaction. Specifically, this procedure is scheduled as follows:
@@
To minimize inference costs, we train a question type classifier on low-resource data and select $E$ based on the predicted type, as further discussed in Appendix A.2.
@@ 3.5 Solutions for Complex Questions
Interactive-KBQA fundamentally involves guiding LLMs through reasoning by annotating an interactive inference process. Thus, for different types of complex questions, it is crucial to identify patterns, design interaction modes, and label high-quality examples. This section outlines several representative types of complex questions to elucidate the design rationale behind this paper.
For multi-hop questions, our focus at each step is on specific predicates rather than on concrete entities. For instance, as depicted in Figure 2, there is no need to locate the specific television programs; expressing the graph pattern within SPARQL is sufficient to allow the tool to handle the ranking part of the task. In the case of CVT structures in Freebase, we explicitly describe the reasoning process upon encountering. Moreover, we break down the star-shaped CVT structure into multiple one-hop relations and treat them as such. For instance, the predicate of “the president of” is represented as “government.governmental_jurisdiction.governing_officials $\rightarrow$ government.government_position_held.office _holder”. When querying the qualifier structure in Wikidata, which modifies predicates, we devise a specialized SPARQL query pattern and provide a detailed thought process as well. Additional examples are provided in the Appendix A.7.
@@ 3.6 Human-Machine Collaborative Annotation
Despite some closed-source LLMs possessing strong reasoning abilities, their utility is limited when their outputs are incorrect and hard to correct. Moreover, existing KBQA datasets primarily annotate the question-logical form pair, omitting detailed reasoning steps. To counter this, we have devised a cooperative human-machine annotation strategy, simplifying data annotation through our interactive KBQA approach. This method entails iteratively modifying LLM outputs at each turn. We manually annotated 50 samples per question type from the training set, creating a valuable low-resource dataset with reasoning process. Moreover, fine-tuning open-source LLMs with this dataset resulted in noticeable performance improvements in low-resource settings, surpassing previous achievements.
@@ 4.1 Dataset & Preprocessing
WebQuestionsSP (WebQSP) Yih et al. (2016) and ComplexWebQuestions 1.1 (CWQ) Talmor and Berant (2018) are extensively utilized in KBQA research and comprise natural language questions alongside their corresponding SPARQL queries based on Freebase. Following Chen et al. (2019), we classify WebQSP questions into 1-hop and 2-hop categories based on the length of the inferential relation chain, i.e., the path from the topic entity to the answer node. CWQ extends WebQSP and encompasses four types of complex questions: conjunction (Conj), composition (Compo), comparative (Compa), and superlative (Super).
KQA Pro Cao et al. (2022) is a large-scale dataset for complex question answering over a dense set of Wikidata. It features nine types of complex questions, which include Count (Ct), QueryAttr (QA), QueryAttrQualifier (QAQ), QueryName (QN), QueryRelation (QR), QueryRelationQualifier (QRQ), SelectAmong (SA), SelectBetween (SB), and Verify (Vf).
MetaQA Zhang et al. (2018) is built upon a Movie Knowledge Base (KB) and includes three sets of question-answer pairs: 1-hop, 2-hop, and 3-hop. We convert the Movie KB into RDF triples, facilitating querying via SPARQL.
Due to the prohibitive costs associated with utilizing OpenAI, we employed a uniform sampling method across each type to construct a smaller test dataset. For each database (DB), we sampled 900 instances, distributing these as 150, 150, 100, and 300 instances per type for WebQSP, CWQ, KQA Pro, and MetaQA, respectively. The statistical details of the test dataset and our annotated (Anno) dataset are presented in Table 1. Our sampled dataset presents a balanced composition, enhancing the assessment of models’ capabilities in handling complex questions.

### v2

@@ Abstract
This study explores the realm of knowledge base question answering (KBQA). KBQA is considered a challenging task, particularly in parsing intricate questions into executable logical forms. Traditional semantic parsing (SP)-based methods require extensive data annotations, which result in significant costs. Recently, the advent of few-shot in-context learning, powered by large language models (LLMs), has showcased promising capabilities. However, fully leveraging LLMs to parse questions into logical forms in low-resource scenarios poses a substantial challenge. To tackle these hurdles, we introduce Interactive-KBQA, a framework designed to generate logical forms through direct interaction with knowledge bases (KBs). Within this framework, we have developed three generic APIs for KB interaction. For each category of complex question, we devised exemplars to guide LLMs through the reasoning processes. Our method achieves competitive results on the WebQuestionsSP, ComplexWebQuestions, KQA Pro, and MetaQA datasets with a minimal number of examples (shots). Importantly, our approach supports manual intervention, allowing for the iterative refinement of LLM outputs. By annotating a dataset with step-wise reasoning processes, we showcase our model’s adaptability and highlight its potential for contributing significant enhancements to the field. [^1] Code and data are available at: https://github.com/JimXiongGM/Interactive-KBQA
@@
Guanming Xiong Peking University Beijing, China gm_xiong@pku.edu.cn                        Junwei Baothanks: Corresponding author. Beijing, China baojunwei001@gmail.com                        Wen Zhao Peking University Beijing, China zhaowen@pku.edu.cn
@@ 1 Introduction
Knowledge base question answering (KBQA) is an increasingly significant research area that leverages structured knowledge bases (KBs) to provide precise answers to natural language (NL) questions. A KBQA system aims to harness the vast, structured information residing in KBs, such as Freebase (Bollacker et al., 2008), Wikidata (Vrandečić and Krötzsch, 2014), or domain-specific databases like the Movie KB (Zhang et al., 2018). Accurately interpreting and responding to user inquiries with data from knowledge bases has potent applications across numerous sectors, making it a focal point of both academic research and industrial innovation.
@@
Underutilization of large language models (LLMs). Despite the demonstrated capabilities of LLMs in reasoning and few-shot learning, existing KBQA approaches have not fully leveraged these strengths. Most current methods either use LLMs as classifiers to identify predicates (Gu et al., 2023; Sun et al., 2024) or as mimics to generate possible logical forms or questions (Li et al., 2023a, b). There remains a significant opportunity to more effectively leverage LLMs in a few-shot setting to enhance the accuracy and interpretability of KBQA systems.
To address these challenges, we introduce the Interactive-KBQA framework, which combines the reasoning power of LLMs with three tools to interact with KBs. By conceptualizing the LLM as an agent and the KB as the environment, Interactive-KBQA facilitates an iterative, dialogue-based problem-solving process. Figure 1 illustrates the overall process. The process adheres to the thought-action paradigm. Given a complex query, the LLM is required to think and provide an action to interact with the KB through a set of specific tools. These tools return the execution results as observations. Specifically, we designed a set of tools that support heterogeneous databases (i.e., Freebase, Wikidata, and a Movie KB) with a unified interaction logic. We categorized complex questions and provided only two annotated exemplars with complete interactive processes for each type, serving as in-context learning demonstrations to prompt the LLM to complete the task. Furthermore, the method introduced in this study allowed for manual intervention. Consequently, we manually annotated a small dataset with a detailed reasoning process, thereby creating a low-resource dataset. Finally, we fine-tuned open-source LLMs on this dataset. The experiments conducted have shown that this method is effective in low-resource contexts. We have released this high-quality dataset to further contribute to NLP research in the community.
@@
  • Design a unified SPARQL-based toolset and interaction logic that efficiently address a wide array of complex queries.
@@ 2 Related Work
Knowledge base question answering (KBQA) methods can be broadly classified into two distinct categories: information retrieval (IR)-based methods and semantic parsing (SP)-based methods. These approaches address the challenge of effectively mapping natural language (NL) queries to the structured formats of knowledge bases (KBs).
@@
SP-based methods parse questions into executable logical forms and perform queries against the KB to retrieve answers. Initial works translated questions into intermediate logical forms before execution. Yih et al. (2015) defined a query graph that resembles subgraphs of the KB and can be directly mapped to a logical form. Hu et al. (2018); Lan and Jiang (2020) extended query graphs to include aggregation operators, and Chen et al. (2021b) enhanced candidate query generation by leveraging predictions of query structure.
In recent years, generative models have increasingly been used to directly generate logical forms. Chen et al. (2021a) designed a retriever to fetch relevant KB items and utilized a Grammar-based Decoder based on LSTM for generating S-expressions (Gu et al., 2021). Meanwhile, Das et al. (2021) employed case-based reasoning for the KBQA task, using Big Bird (Zaheer et al., 2020) to generate logical forms by retrieving relevant cases. Notably, works such as Ye et al. (2022), Shu et al. (2022), Zhang et al. (2023), and Yu et al. (2023) involved initially retrieving basic elements from the KB, such as entities, relations, subgraphs, or texts as supplementary information, before directly generating logical forms using the T5 model (Raffel et al., 2020). However, this approach requires substantial amounts of training data and suffers from a lack of transparency in the reasoning process.
Large language models (LLMs) for KBQA leverage the potent few-shot learning capabilities inherent in LLMs. Recent studies have shown that LLMs can significantly enhance reasoning over KBs. By capitalizing on the powerful few-shot learning capabilities, these methods have demonstrated marked improvements in both the accuracy and efficiency of information retrieval from KBs.
In the context of few-shot scenarios, Gu et al. (2023) proposed that LLMs should prioritize evaluating the plausibility of agent plans over directly generating answers. Concurrently, Li et al. (2023a) advocated for the generation of logical forms as initial drafts, which are subsequently refined into executable queries using the KB. From the agent-environment perspective, Jiang et al. (2023b) developed two specialized interfaces for accessing the KB, while Gu et al. (2024) and Liu et al. (2024) designed seven tools to facilitate this process. Sun et al. (2024) introduced a novel approach that enables LLMs to iteratively employ beam search reasoning on a KB. Furthermore, Zong et al. (2024) proposed assigning three distinct roles to the agent for addressing different KBQA subtasks. Jiang et al. (2024) constructed an instruction dataset based on existing KBQA datasets. However, despite employing LLMs, these methods lack the exploration of more complex questions.
@@ 3.1 Problem Formulation
This study investigates a semantic parsing (SP) method for knowledge base question answering (KBQA). A knowledge base (KB) is formally represented as $K\in E\times R\times(E\cup L\cup C)$, where $E$ denotes the set of entities, $R$ signifies the set of relations between entities, $C$ represents the set of classes, and $L$ includes the literal values. Given a question $Q$ and a knowledge base $\mathcal{K}$, our goal is to generate an executable SPARQL expression $S$ that aligns with the question. Thus, the task can be formalized as $p(S|Q,\mathcal{K})$.
@@ 3.2 Overview
Recent advancements in large language models (LLMs) have demonstrated remarkable capabilities in few-shot learning and reasoning. However, fully leveraging LLMs to tackle complex KBQA challenges remains an elusive goal. To address this gap, we introduce Interactive-KBQA, an interactive method for KBQA that conceptualizes the LLM as an agent and the KB as an environment. This approach enables semantic parsing and SPARQL generation through dialogic interactions. Specifically, we design a unified interaction logic using three generic tools capable of supporting various complex types of questions (e.g., type constraints, count-based queries) across multiple databases (e.g., Freebase, Wikidata, Movie KB). Figure 3 presents an illustrative example of the interactive process.
@@ 3.3 Tools for Knowledge Base
The SP-based KBQA method necessitates the identification of elements and the appropriate graph patterns. Hence, it is crucial to design tools at an atomic level to ensure their universality. In line with this principle, we introduce the following three tools.
SearchNodes(name): This function searches for nodes in the KB using the given surface name, name. Its primary purpose is entity linking (EL). Consequently, it not only returns the formal name of a node but also provides distinguishing features of the entity, such as its description and type. Importantly, this tool avoids dataset-specific EL techniques in favor of a generic retrieval approach.
SearchGraphPatterns(sparql, semantic): This function aims to identify and rank graph predicates within the KB that are essential, guided by the semantic parameter. The function requires the input sparql, which should consist of a SPARQL query beginning with "SELECT ?e WHERE ". Following this, it performs a query on the one-hop subgraph centered around the entity ?e. This includes both incoming and outgoing edges. Subsequently, it ranks the retrieved triples based on the semantic relevance to the semantic parameter and the predicates in triples. Ultimately, the tool returns the top K triples. This tool is designed to precisely identify subgraphs while discarding irrelevant information, thereby optimizing the context window’s use in LLMs. It supports flexible operations and is specifically optimized for Compound Value Type [^4] A Compound Value Type is a Type within Freebase designed to represent data where each entry is composed of multiple fields. (CVT) in Freebase by flattening a CVT node to multiple single-hop relationships. For example, to find movies featuring Tom Hanks in Freebase, the usage would be: SearchGraphPatterns(’SELECT ?e WHERE {?e type.object.name "Tom Hanks"@en .}’, semantic="play in film"), and return: [(?e, film.actor.film -> film.performance.film, "Nothing in Common"), ...].
@@
Implementation details are provided in Appendix A.2.1, and additional usage instructions are described in Appendix A.2.3.
@@ 3.4 Interactive Process
Given a question $Q$, we first construct a prompt text:
@@
where the instruction text, Inst, consists of tool descriptions, tool usage, and the format of interaction. Inst is database-specific and pre-written. $E$ denotes a set of exemplars, and for each type of question, we manually annotate two complete examples in an interactive format.
In each turn $T$, we let the LLM generate an action based on the Prompt and the history $H$ of the interaction. Specifically, this procedure follows:
@@
To minimize inference costs, we train a question type classifier on low-resource data and select $E$ based on the predicted type, as further discussed in Appendix A.2.1.
@@ 3.5 Solutions for Complex Questions
Interactive-KBQA involves guiding LLMs through reasoning by annotating an interactive inference process. For different types of complex questions, it is crucial to identify patterns, design interaction modes, and label high-quality examples. This section outlines several representative types of complex questions to elucidate the design rationale behind this paper.
For multi-hop questions, our focus at each step is on specific predicates rather than concrete entities. For instance, as depicted in Figure 3, there is no need to locate specific television programs; expressing the graph pattern within SPARQL is sufficient to enable the tool to handle the ranking part of the task. In the case of CVT structures in Freebase, we explicitly describe the reasoning process when encountered. Moreover, we break down the star-shaped CVT structure into multiple one-hop relations and treat them accordingly. For instance, the semantics of the sentence “Tom Hanks plays the role of ‘David Basner’ in the film ‘Nothing in Common’ ” can be represented by two triples: (“Tom Hanks”, film.actor.film $\rightarrow$ film.performance.film, “Nothing in Common”) and (“Tom Hanks”, film.actor.film $\rightarrow$ film.performance.character, “David Basner”). When querying the qualifier structure in Wikidata, which modifies predicates, we devise a specialized SPARQL query pattern and provide a detailed thought process as well. Additional examples are provided in the Appendix A.8.
@@ 3.6 Human-Machine Collaborative Annotation
In realistic scenarios, including examples of all question types within the context is impractical for two primary reasons. First, the associated costs are substantial. Second, as discussed by Su et al. (2023); Zhu et al. (2023), LLMs experience a notable decline in performance when input tokens exceed certain limits. Additionally, Lightman et al. (2023) pointed out that process supervision can enhance a model’s generalization capabilities. Consequently, we propose a human-machine collaborative data annotation method. Enabled by the Interactive-KBQA method introduced in this paper, annotating the reasoning process has become more straightforward.
Specifically, when annotators determine that the action $a_{T}$ is unreasonable, it is manually corrected to $a^{\prime}_{T}$. This adjustment is then incorporated into the message to generate the action $a_{T+1}$. Formally,
$a_{T+1}=\text{LLM}(\{\text{Prompt},\{c_{0},a_{0},o_{0},...,c_{T-1},a^{\prime}_{T},o^{\prime}_{T}\}\})$
We set a breakpoint after each round where the LLM generates thoughts and actions, allowing human evaluators to review and decide on their acceptance. Once accepted, the process continues. If rejected, the annotator revises the generated thoughts and actions before proceeding. For each question type, we randomly sampled 50 data points from the training set and manually annotated them to provide low-resource data.
The core principle of our methodology is to emulate the human data annotation process. Annotators intervene in specific scenarios where the model exhibits hallucinations (such as generating predicates not present in the observation), inconsistencies between Thought and Action, or deviations from the correct answer trajectory (for instance, cases where two consecutive rounds are incorrect).
@@ 4.1 Dataset & Preprocessing
WebQuestionsSP (WebQSP) (Yih et al., 2016) and ComplexWebQuestions 1.1 (CWQ) (Talmor and Berant, 2018) are extensively used in KBQA research. These datasets comprise natural language questions paired with their corresponding SPARQL queries based on Freebase. Following Chen et al. (2019), we classify WebQSP questions into 1-hop and 2-hop categories based on the length of the inferential relation chain, i.e., the path from the topic entity to the answer node. CWQ extends WebQSP by incorporating four types of complex questions: Conjunction (Conj), Composition (Compo), Comparative (Compa), and Superlative (Super).
KQA Pro (Cao et al., 2022) is a large-scale dataset designed for complex question answering over a dense set of Wikidata entries. It features nine types of complex questions, including Count (Ct), Query Attribute (QA), Query Attribute Qualifier (QAQ), Query Name (QN), Query Relation (QR), Query Relation Qualifier (QRQ), Select Among (SA), Select Between (SB), and Verify (Vf).
MetaQA (Zhang et al., 2018) is built upon a Movie Knowledge Base (KB) and includes three sets of question-answer pairs: 1-hop, 2-hop, and 3-hop. We converted the Movie KB into RDF triples to facilitate querying via SPARQL.
Due to the prohibitive costs associated with utilizing OpenAI, we employed a uniform sampling method across each type to construct a smaller test dataset. For each DB, we sampled 900 instances, distributing these as follows: 150 instances each for WebQSP and CWQ, 100 instances for KQA Pro, and 300 instances for MetaQA. The statistical details of the test dataset and our annotated (Anno) dataset are presented in Table 1. Our sampled dataset presents a balanced composition, enhancing the assessment of models’ capabilities in handling complex questions.